OpenStudy (anonymous):

7(3x+2)^2(1-x)^2+(3x+2)(1-x)^3=

OpenStudy (anonymous):

are you supposed to multiply out or factor?

OpenStudy (anonymous):

factor compleatly and simplify

OpenStudy (anonymous):

completely

OpenStudy (anonymous):

ok then ignore my answer to your last problem. i thought it meant multiply

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this one factors as \[5 (x-1)^2 (3 x+2) (4 x+3)\]

OpenStudy (anonymous):

the 1st one

OpenStudy (anonymous):

each term has a common factor of \[(3x+2)(1-x)^2\]

OpenStudy (anonymous):

you need an answer to your first question as well?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[(4 x-3) (60 x-23)\]

OpenStudy (anonymous):

that is the answer to the one i multiplied out earlier. there it is in factored form

OpenStudy (anonymous):

i have to work the whole thing out and i really am not good at this

OpenStudy (anonymous):

ok which one would you like to do?

OpenStudy (anonymous):

the one that started with 5 thanks

OpenStudy (anonymous):

let me go find it hold on

OpenStudy (anonymous):

\[5(3-4x)^2-8(3-4x)(5x-1)=\]

OpenStudy (anonymous):

ok this one is easier than the second one. each term has a common factor. it is \[3-4x\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so before we do this, lets imagine we had \[5a^2-8ab\] and we wanted to factor

OpenStudy (anonymous):

i see a common factor of "a" in each term so i "factor it out" and write \[a(5a-8b)\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

we are going to do exactly the same thing, but with \[a=3-4x \text{ and } b=5x-1\]

OpenStudy (anonymous):

so we write \[(3-4x)(5(3-4x)-8(5x-1))\]

OpenStudy (anonymous):

so we write \[(3-4x)(5(3-4x)-8(5x-1))\]

OpenStudy (anonymous):

i am guessing that is the confusing step, so let me know if it is not clear

OpenStudy (anonymous):

but the line above looks just like \[a(5a-8b)\] as i wrote before, but with the replacements for a and for b

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes as in "yes it is confusing"? or yes as in "ok i get it more or less"

OpenStudy (anonymous):

where is the 2nd part of 3-4x

OpenStudy (anonymous):

confusing

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

we start with \[5(3-4x)^2-8(3-4x)(5x-1)=\]

OpenStudy (anonymous):

each term has a common factor. it is \[(3x-4)\] that much ok?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

ok so what is it being multiplied by in the second term? it is being multiplied by \[-8(5x-1)\] because you see \[-8(3x-4)(5x-1)\]

OpenStudy (anonymous):

and in the first term you have \[5(3x-4)=5(3x-4)(3x-4)\]

OpenStudy (anonymous):

so in the first term it is being multiplied by \[5(3x-4)\]

OpenStudy (anonymous):